OpenStudy (anonymous):

please help how do u do this!! Evaluate each expression and match with the correct value. P(7,5) P(8,2) C(5,2) C(12,7) C(12,4) * C(8,3) Choose... 2720 27,720 10 792 2520 56

OpenStudy (anonymous):

are you working out of the Glencoe Algebra II book

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I was just wondering because these questions are in my book too

OpenStudy (anonymous):

this one I am not to good at

OpenStudy (anonymous):

can u try i have no idea

OpenStudy (anonymous):

are you using a book or is it just worksheets

OpenStudy (anonymous):

worksheets

OpenStudy (anonymous):

does it give you and idea of what section it might be from

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

okay is this different questions or all one

OpenStudy (anonymous):

all one

OpenStudy (anonymous):

so choose at the problem is the answers

OpenStudy (anonymous):

because if that is the case then this would be what you do nPr(8,2) 56 = 8!/(8-2)! = 8*7*6*5*4*3*2*1/6*5*4*3*2*1 nPr(7,5) 2520 = 7!/(7-5)! = 7*6*5*4*3*2*1/2*1 nCr(5,2) 10 = 5!/(5-2)!*2! = 5*4*3*2*1/(3*2*1)(2*1) nCr(12,7) 792 nCr(12,4)*nCr(8,3) 27720

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

okay I will redo P(8,2) P(8,2) 8!/(8-2)! = 8*7*6*5*4*3*2*1/6*5*4*3*2*1

OpenStudy (anonymous):

which equals 56